OpenStudy (anonymous):

find the solutions to the equation below. Check all that apply. 4x^2-4x-3=0 A.X=-1/2 B.x=-2 C.x=1/2 D.x=2 E.x=3/2 F. X=3

OpenStudy (anonymous):

Can you do this one on your own? Or do you want me to start it?

OpenStudy (anonymous):

answers have to + huh

OpenStudy (anonymous):

let me try and just check my answer

OpenStudy (anonymous):

No necessarily.... When you answer is like the one below, you have a positive AND a negative\[ax^2 - x - b^2\]

OpenStudy (anonymous):

got it A nd E

OpenStudy (anonymous):

YUPPERS

OpenStudy (anonymous):

\[4x^2-4x-3=0 \]\[(2 x-3) (2 x+1) = 0 \]\[x = -\frac{ 1 }{ 2 }, \frac{ 3 }{ 2 }\]